Daily Enterprises is purchasing a $9.9 million machine. It will cost $52,000 to transport and install the machine. The machine has a depreciable life of five years and will have no salvage value. The machine will generate incremental revenues of $3.9 million per year along with incremental costs of $1.4 million per year. If​ Daily's marginal tax rate is 35%​, what are the incremental earnings​ (net income) associated with the new​ machine? The annual incremental earnings are. (Round to the nearest​ dollar.)
Answer #1

Incremental revenue = $3.9 million

Incremental cost = $ 1.4 million

Depreciation = Cost / useful life

= $9.9 + $52000 / 5

= 9.952/5 = 1.9904

Incremental earning = (revenue - cost - depreciation ) ( 1 - tax rate)

= ( 3.9 - 1.4 - 1.9904 ) ( 1 - 0.35)

= 0.5096 * 0.65

= $0.33124 million